OpenStudy (anonymous):

(sin θ + cos θ) ^ 2 / (1+2sinθcosθ) = ??

OpenStudy (anonymous):

Answer Choices: -sec^ 2 θ 1 0 1 - sin θ

OpenStudy (anonymous):

sin2 + cos2 = 1 so (sin + cos)^2 = sin2 + 2sin cos + cos2 =1 + 2sin cos then the answer is 1

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I love identities

OpenStudy (anonymous):

I hate them -_-